OpenStudy (anonymous):

Solve the following system of equations. x - 2y = 5 2x - 4y = 10 Infinitely Many Solutions No Solutions (0, 0) (5, 10)

OpenStudy (anonymous):

Isolate both y's on your two systems and the use y1 = y2

OpenStudy (anonymous):

\[ \ \ x-2y=5\\ \ \ 2x-4y=10\\ \text{First step is to eliminate a variable, so multiply the first by -2.}\\ \ \ -2[x-2y]=-2[5]\\ \ \ -2x+4y=-10\\ \text{Combine the equations.}\\ \ \ 0=0\\ \text{Since this is an identity, these lines are identical.}\\ \text{Therefore there are infinitely many solutions.} \]

OpenStudy (anonymous):

infinitely many solutins since both eqn s are the same!!!!!!!!!!!!! 2x-4y=10 2(x-2y)=10 x-2y=5